At the temperature celsius will 19.4 g of molecular oxgem, O 2 expert a pressure of 1820 mm hg in a 5.12 l cylinder worked out
1
Expert's answer
2020-04-19T07:09:58-0400

pV=(m/Mr) *R*T;

T=p*V*Mr/(m*R)=242.64*5.12*32/19.4*8.314=246.47 K =-26.53 oC